OpenStudy (anonymous):

A 15 ft. diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 3 ft. longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by the brace.

OpenStudy (campbell_st):

 draw a picture then use pythagoras' theorem for the given information \[h^2 + (h +3)^2 = 15^2\] expand and collect like terms \[2h^2 +6h - -216 = 0\] solve the quadratic and only use positive values for h.... as distance can only be positive